# Efficient Large-Scale Quantum Optimization via Counterdiabatic Ansatz

### v1

@@ III.2 Effective Dimension and Excitation Suppression
III.3 Instance-Sequential Training (IST)
@@
To fully leverage the advantage afforded by such phenomenon of parameter concentration in DC-QAOA(NC), we propose the IST method, that can make the training process of DC-QAOA(NC) more efficient in using quantum resources. Under the IST method, for a graph with $N$ vertices, we randomly remove $N-k$ vertices (and the edges associated with them) while keeping the graph connected, i.e. the graph should not be separated to two. Firstly, we train DC-QAOA(NC) on the remaining $k$ vertices to get optimal parameters $(\boldsymbol{\alpha}_{k},\boldsymbol{\beta}_{k},\boldsymbol{\gamma}_{k})$. Then, one vertex is added back to the graph (and the edges associated with it) and DC-QAOA(NC) is trained again using $(\boldsymbol{\alpha}_{k},\boldsymbol{\beta}_{k},\boldsymbol{\gamma}_{k})$ as the starting point to obtain $(\boldsymbol{\alpha}_{k+1},\boldsymbol{\beta}_{k+1},\boldsymbol{\gamma}_{k+1})$. We then add one more vertex back, and the above steps are repeated until $(\boldsymbol{\alpha}_{N},\boldsymbol{\beta}_{N},\boldsymbol{\gamma}_{N})$ is obtained.
@@
To demonstrate the effectiveness of the IST method, we compare it to the traditional training method while fixing the number of total training iterations the same. In this comparison, the IST method uses much fewer quantum resources because it does not have to access all $N$ qubits in the quantum processing unit in every iteration. We benchmark the advantage of performance by the proportion of approximate ratio $R_{\mathrm{IST}}$ of the IST method relative to the one for traditional method $R_{\mathrm{traditional}}$. In Fig. 8, the performance of DC-QAOA(NC) trained using IST method is nearly the same with the one trained using traditional method using much less quantum resources. Sometimes, for $N=9$ and 13 DC-QAOA(NC) trained using IST works even better than the one trained using traditional method. However, for QAOA, the IST method fails due to the lack of parameter concentration. We conclude that when the number of training iterations are held same for both methods, IST method, on average, takes less time to run a quantum devices and use fewer qubits to construct the circuits. This is a favorable characteristic especially when the access to the device is limited by certain quota, e.g. cloud-based ones.
@@
Figure 8: A comparative analysis between IST and traditional training methods in DC-QAOA(NC) and QAOA. A dashed line is drawn at $R_{\mathrm{IST}}/R_{\mathrm{traditional}}=1$ as a guide. IST not only matches but sometimes surpasses traditional methods in achieving optimal expectation values, demonstrating superior efficiency with significantly fewer quantum resources because parameter concentration.
IV CONCLUSION
@@
In this work, we have used the MaxCut problem as an example to throughly examine the comparison between DC-QAOA and QAOA. Such a comparison has been previously done with the number of layers held equal and with a conclusion that DC-QAOA is superior to QAOA for all sizes of the problem. However, we demonstrate that this statement must be carefully refined according to the actual circuit complexity afforded by different algorithms, which includes both the number of layers and number of CNOT gates in each layer. The cd-terms introduced in DC-QAOA reduces the critical depth of the algorithm but also introduces additional CNOT gates, so the overall circuit complexity is not necessarily lower than QAOA without cd-terms. We have found that for MaxCut problem with $N<16$, QAOA is more advantageous than DC-QAOA(NC), while for larger problems with $N>16$, DC-QAOA(NC) is superior. This refinement of the comparison between the two algorithms has not been previously noted in the literature. Although both YY and Y types of cd-terms enhance the performance of QAOA, the magnitude of improvement is not as significant as that achieved by DC-QAOA(NC).
@@
Furthermore, we calculated the effective dimensions of the algorithms and have shown that the superior performance of DC-QAOA(NC) in larger problems is rooted in its strong ability to suppress unwanted excitation and to add the effective dimensions. We have also noted that the optimal parameters found by DC-QAOA(NC) is more concentrated than those from QAOA. Based on this finding, we introduced an IST training method, which achieves similar (and sometimes even better) results as compared to traditional trining methods, while consuming much fewer quantum resources. Our results should be valuable to recent efforts in utilizing VQA to solve combinatorial optimization problems for large-scale computational tasks, thereby advancing the study of quantum optimization.
@@ ACKNOWLEDGEMENT
This work is supported by the National Natural Science Foundation of China (Grant No. 11874312, 12474489), the Research Grants Council of Hong Kong (CityU 11304920), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2303007), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302300).
APPENDIX

### v2

@@ III.3 Effective Dimension and Excitation Suppression
III.4 Instance-Sequential Training (IST)
@@
To fully leverage the advantage afforded by such phenomenon of parameter concentration in DC-QAOA(NC), we propose the IST method, that can make the training process of DC-QAOA(NC) more efficient in using quantum resources. Under the IST method, for a graph with $N$ vertices, we randomly remove $N-k$ vertices (and the edges associated with them) while keeping the graph connected, i.e. the graph should not be separated to two. Firstly, we train DC-QAOA(NC) on the remaining $k$ vertices to get optimal parameters $(\boldsymbol{\alpha}_{k},\boldsymbol{\beta}_{k},\boldsymbol{\gamma}_{k})$. Then, one vertex is added back to the graph (and the edges associated with it) and DC-QAOA(NC) is trained again using $(\boldsymbol{\alpha}_{k},\boldsymbol{\beta}_{k},\boldsymbol{\gamma}_{k})$ as the starting point to obtain $(\boldsymbol{\alpha}_{k+1},\boldsymbol{\beta}_{k+1},\boldsymbol{\gamma}_{k+1})$. We then add one more vertex back, and the above steps are repeated until $(\boldsymbol{\alpha}_{N},\boldsymbol{\beta}_{N},\boldsymbol{\gamma}_{N})$ is obtained. The detailed procedure of the IST method is outlined in Algorithm 1, and an illustrative flow chart is presented in Fig. 9.
Algorithm 1 Instance-Sequential Training (IST) Method
0:  Graph $G$ with $N$ vertices
1:  Initialize an empty set of parameters $\Theta$
2:  Randomly remove $N-k$ vertices to form a reduced graph $G_{k}$, ensuring connectivity
3:  Train the DC-QAOA circuit on $G_{k}$ to find optimal parameters $\Theta_{k}$
4:  Set $\Theta=\Theta_{k}$
5:  for $i=k+1$ to $N$ do
6:     Add one vertex back to $G$ to form $G_{i}$
7:     Train DC-QAOA on $G_{i}$ using $\Theta$ as the initial parameters
8:     Update $\Theta$ with the new optimal parameters
9:  end for
10:  return  Final parameters $\Theta$
Figure 9: An illustrative flow chart of IST algorithm. In the illustration, we use a six-node graph as an example. The original graph is shown above ”start”, after initialization of $\Theta$ with $\Theta_{0}$, two nodes are removed from the graph, together with all the edges connected to these nodes, shown by red dashed line. In the loop process, one node is added back to the graph to conduct another round of training, together with its edges with all the existing nodes, also shown by red dashed line.
@@
To demonstrate the effectiveness of the IST method, we compare it to the traditional training method while fixing the number of total training iterations the same. In this comparison, the IST method uses much fewer quantum resources because it does not have to access all $N$ qubits in the quantum processing unit in every iteration. We benchmark the advantage of performance by the proportion of approximate ratio $R_{\mathrm{IST}}$ of the IST method relative to the one for traditional method $R_{\mathrm{traditional}}$. In Fig. 11, the performance of DC-QAOA(NC) trained using IST method is nearly the same with the one trained using traditional method using much less quantum resources. Sometimes, for $N=9$ and 13 DC-QAOA(NC) trained using IST works even better than the one trained using traditional method. However, for QAOA, the IST method fails due to the lack of parameter concentration. We conclude that when the number of training iterations are held the same for both methods, IST method, on average, takes less time to run a quantum device and uses fewer qubits to construct the circuits. This is a favorable characteristic especially when the access to the device is limited by certain quota, e.g. cloud-based ones.
@@
Figure 11: A comparative analysis between IST and traditional training methods in DC-QAOA(NC) and QAOA. IST not only matches but sometimes surpasses traditional methods in achieving optimal expectation values, demonstrating superior efficiency with significantly fewer quantum resources because parameter concentration.
IV CONCLUSION and Outlook
@@
In this work, we have used the MaxCut problem as an example to thoroughly examine the comparison between DC-QAOA and QAOA. Such a comparison has been previously done with the number of layers held equal and with a conclusion that DC-QAOA is superior to QAOA for all sizes of the problem. However, we demonstrate that this statement must be carefully refined according to the actual circuit complexity afforded by different algorithms, which includes both the number of layers and number of CNOT gates in each layer. The cd-terms introduced in DC-QAOA reduces the critical depth of the algorithm but also introduces additional CNOT gates, so the overall circuit complexity is not necessarily lower than QAOA without cd-terms. We have found that for MaxCut problem with $N<16$, QAOA is more advantageous than DC-QAOA(NC), while for larger problems with $N>16$, DC-QAOA(NC) is superior. This refinement of the comparison between the two algorithms has not been previously noted in the literature. Although both YY and Y types of cd-terms enhance the performance of QAOA, the magnitude of improvement is not as significant as that achieved by DC-QAOA(NC).
To further validate the practicality of DC-QAOA, we benchmark one-layer DC-QAOA(NC) against three-layer QAOA, ensuring they use the same number of CNOT gates. In noiseless simulations, we find that DC-QAOA(NC) starts outperforming QAOA when the problem size exceeds $N=16$. More importantly, in a noisy environment simulated using a depolarization model with realistic error rates in NISQ devices, DC-QAOA(NC) achieves an advantage at an even smaller problem size ($N\approx 11$). This result highlights the robustness of DC-QAOA(NC) against noise, making it a promising approach for near-term quantum hardware.
@@
Furthermore, we calculated the effective dimensions of the algorithms and have shown that the superior performance of DC-QAOA(NC) in larger problems is rooted in its strong ability to suppress unwanted excitation and to add the effective dimensions. We have also noted that the optimal parameters found by DC-QAOA(NC) are more concentrated than those from QAOA. Based on this finding, we introduced an IST training method, which achieves similar (and sometimes even better) results as compared to traditional training methods, while consuming much fewer quantum resources. Our results should be valuable to recent efforts in utilizing VQA to solve combinatorial optimization problems for large-scale computational tasks, thereby advancing the study of quantum optimization.
While our study focuses on comparing DC-QAOA with standard QAOA, it is important to acknowledge that numerous other QAOA variants have been explored, each designed to enhance specific aspects of the algorithm. Broadly, these variants can be classified into three main categories. The first category aims to increase the expressivity and flexibility of the ansatz by introducing more variational parameters, as seen in Expressive QAOA (XQAOA) Vijendran et al. (2024) and Multi-Angle QAOA (MA-QAOA) Herrman et al. (2022); Gaidai and Herrman (2024). Since standard QAOA can be viewed as a subset of these approaches, they ideally perform at least as well as standard QAOA in the worst case. The same motivation can be extended to DC-QAOA, as adding more parameters could potentially enhance its optimization capability. However, this comes at the cost of increased sampling overhead, as estimating gradients for a higher number of parameters demands more measurements, exposing the algorithm to greater shot noise. This issue is particularly significant in the NISQ era, where hardware constraints limit measurement precision. Furthermore, overparameterization alters the loss landscape of the cost function, potentially leading to trainability issues such as barren plateaus. Future studies should carefully examine whether adding more parameters to DC-QAOA (particularly DC-QAOA(NC)) yields a genuine performance improvement or simply exacerbates these challenges.
The second category of QAOA variants focuses on redesigning the mixing operator instead of relying on the problem Hamiltonian $H_{f}$, as in the case of Grover Mixer QAOA (GM-QAOA) Bridi and Marquezino (2024); Bärtschi and Eidenbenz (2020). This poses a fundamental challenge for DC-QAOA because it originates from counterdiabatic driving, which depends on constructing an appropriate gauge potential to suppress excitations. When a different mixing Hamiltonian is introduced, designing a corresponding counterdiabatic term becomes nontrivial, potentially requiring new approaches to ensure diabatic transitions are effectively minimized. Exploring how counterdiabatic strategies can be adapted to non-standard mixing operators presents an interesting direction for future research.
A third category of variants seeks to reduce the quantum computational burden by leveraging classical preprocessing techniques that exploit problem structure, as seen in Recursive QAOA (RQAOA) Bravyi et al. (2020) and ST-Spanning Tree QAOA (ST-QAOA) Wurtz and Love (2021). These methods apply classically solvable criteria to simplify the problem before executing the quantum algorithm. Unlike the previous categories, these techniques can be seamlessly integrated into DC-QAOA without fundamentally altering its counterdiabatic framework. Future work could explore how such classical-quantum hybrid strategies can further enhance the practical implementation of DC-QAOA, particularly in scaling to larger problem instances while mitigating quantum resource constraints.
Overall, our findings suggest that DC-QAOA provides a meaningful alternative to standard QAOA, particularly when considering circuit complexity in the NISQ era. However, further research is needed to explore its applicability across broader problem classes, its potential integration with other QAOA variants, and its behavior under realistic quantum hardware noise. A deeper understanding of its trainability, parameter concentration properties, and adaptation to alternative ansatz will be essential in determining its viability for large-scale quantum optimization in the near future.
@@ ACKNOWLEDGEMENT
This work is supported by the National Natural Science Foundation of China (Grant Nos. 11874312 and 12474489), the Research Grants Council of Hong Kong (CityU 11304920), Shenzhen Fundamental Research Program (Grant No. JCYJ20240813153139050), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2403001), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302300).
Appendix A Proof of equivalence between Max-k-SAT and MaxCut
Max-k-SAT and MaxCut are two fundamental problems in combinatorial optimization and computational complexity. Both are known to be NP-hard and widely studied in theoretical computer science. In this appendix, we present a detailed polynomial-time reduction between Max-k-SAT and MaxCut in both directions, thereby demonstrating their computational equivalence. While such reductions are standard (see, e.g., Garey and Johnson (1990); Goemans and Williamson (1995b); Li et al. (2007)), we include them here for completeness and pedagogical clarity, particularly to motivate why the logistic saturation conjecture, initially formulated for Max-2-SAT, can be extended to MaxCut.
A.0.1 Max-k-SAT
Let $\varphi$ be a Boolean formula in Conjunctive Normal Form (CNF), which consists of $m$ clauses, each containing at most $k$ literals. A literal is a variable $x_{i}$ or its negation $\neg x_{i}$, where $x_{i}\in\{x_{1},x_{2},\dots,x_{n}\}$ is a Boolean variable.
We define the Max-k-SAT problem as follows:
  1. 1.
    Input:
    • A set of Boolean variables $\{x_{1},x_{2},\dots,x_{n}\}$.
    • A Boolean formula $\varphi=C_{1}\land C_{2}\land\cdots\land C_{m}$, where each clause $C_{i}$ is a disjunction of at most $k$ literals, i.e.,
      $C_{i}=\ell_{i1}\vee\ell_{i2}\vee\cdots\vee\ell_{ik}$
      where each $\ell_{ij}$ is a literal which is either a variable or its negation (for example, $x$ or $\neg x$). In this way, $\ell_{ij}\in\{x_{1},x_{2},\dots,x_{n},\neg x_{1},\neg x_{2},\dots,\neg x_{n}\}$.
  2. 2.
    Goal: Find a truth assignment $\mathbf{A}=(a_{1},a_{2},\dots,a_{n})$ where $a_{i}\in\{\text{True},\text{False}\}$ for every Boolean variable, such that the number of satisfied clauses is maximized.
  3. 3.
    Satisfaction of a Clause: A clause $C_{i}=\ell_{i1}\vee\ell_{i2}\vee\cdots\vee\ell_{ik}$ is said to be satisfied if at least one of its literals evaluates to true under the truth assignment $\mathbf{A}$. Formally, for a clause $C_{i}$, it is satisfied if there exists at least one $j\in\{1,2,\dots,k\}$ such that:
    $\text{Evaluate}(\ell_{ij},\mathbf{A})=\text{True}$
    where $\text{Evaluate}(\ell_{ij},\mathbf{A})=\text{True}$ if $\ell_{ij}$ evaluates to true under the assignment $\mathbf{A}$, and False otherwise.
  4. 4.
    Objective: Maximize the number of satisfied clauses, i.e., the goal is to find the assignment $\mathbf{A}$ such that the number of satisfied clauses $\text{Satisfy}(\mathbf{A})$, defined as:
    $\text{Satisfy}(\mathbf{A})=\left|\left\{i\mid C_{i}\text{ is satisfied by }\mathbf{A}\right\}\right|$
    is maximized.
  5. 5.
    Maximization: The problem is to find the truth assignment $\mathbf{A}$ that maximizes $\text{Satisfy}(\mathbf{A})$. Mathematically:
    $\mathbf{A^{*}}=\arg\max_{\mathbf{A}}\text{Satisfy}(\mathbf{A})$
    where $\mathbf{A^{*}}$ is the truth assignment that maximizes the number of satisfied clauses.
In summary, Max-k-SAT is a maximization problem where the aim is to optimize the truth assignment to variables such that the total number of clauses that evaluate to true is as large as possible. One example of Max-3-SAT is shown in TABLE 4 where for each of the $82^{3}=$ 8 possible truth assignments, we indicate which clauses are satisfied and count the total. The goal in Max-3-SAT is to find the assignment that satisfies the most clauses. In this example, 5 assignments satisfy all three clauses, achieving the maximum value.
$x_{1}$ $x_{2}$ $x_{3}$ Clause 1 Clause 2 Clause 3 # Satisfied
T T T Satisfied Satisfied Satisfied 3
T T F Satisfied Satisfied Satisfied 3
T F T Satisfied Not Satisfied Satisfied 2
T F F Satisfied Satisfied Satisfied 3
F T T Satisfied Satisfied Satisfied 3
F T F Satisfied Satisfied Not Satisfied 2
F F T Satisfied Satisfied Satisfied 3
F F F Not Satisfied Satisfied Satisfied 2
Table 4: Truth table for a Max-3-SAT instance with clauses $C_{1}=x_{1}\lor x_{2}\lor x_{3}$, $C_{2}=\lnot x_{1}\lor x_{2}\lor\lnot x_{3}$, and $C_{3}=x_{1}\lor\lnot x_{2}\lor x_{3}$. The final column shows how many clauses are satisfied by each assignment. The maximum is highlighted.
A.0.2 Max-Cut
The Max-Cut problem is a well-known optimization problem in graph theory. Given a graph, the goal is to partition the vertices of the graph into two disjoint sets such that the number of the edges crossing the partition is maximized.
We define the Max-Cut problem as follows:
  1. 1.
    Input:
    • A graph $G=(V,E)$, where:
      • $V=\{v_{1},v_{2},\dots,v_{n}\}$ is the set of vertices, with $n$ vertices.
      • $E=\{e_{1},e_{2},\dots,e_{m}\}$ is the set of edges, where each edge $e_{i}=(v_{a},v_{b})$ connects two vertices $v_{a},v_{b}\in V$.
    • A partition of the vertices into two sets $V_{1}\subseteq V$ and $V_{2}=V\setminus V_{1}$, where $V_{2}$ is the complement of $V_{1}$, meaning $V_{1}\cup V_{2}=V$ and $V_{1}\cap V_{2}=\emptyset$.
  2. 2.
    Goal: Find a partition $(V_{1},V_{2})$ of the vertex set $V$ such that the number of edges between $V_{1}$ and $V_{2}$ is maximized. That is, we want to maximize:
    $\text{Cut}(V_{1},V_{2})=\left|\left\{e\in E\mid e\text{ connects a vertex in }V_{1}\text{ to a vertex in }V_{2}\right\}\right|$
    where the cut size is the number of edges that have one endpoint in $V_{1}$ and the other in $V_{2}$.
  3. 3.
    Objective: Maximize the value of the cut, which is the number of edges crossing the partition $(V_{1},V_{2})$:
[…]
    where $V_{1}\subseteq V$ is the set of vertices in one partition, and $V_{2}$ is the complement of $V_{1}$ in the partition.
  4. 4.
    Maximization: The problem is to find the partition $(V_{1}^{*},V_{2}^{*})$ that maximizes the cut size. Mathematically:
    $(V_{1}^{*},V_{2}^{*})=\arg\max_{(V_{1},V_{2})}\text{Cut}(V_{1},V_{2})$
    where $(V_{1}^{*},V_{2}^{*})$ is the partition that maximizes the number of edges crossing the cut.
A.1 Reduction from Max-k-SAT to MaxCut
Theorem 1.
Any instance of Max-k-SAT can be transformed into an equivalent instance of MaxCut in polynomial time, preserving the optimization structure and solution correspondence.
Proof.
Construction: Given a Max-k-SAT instance with variables $x_{1},\dots,x_{n}$ and clauses $C_{1},\dots,C_{m}$, construct a weighted graph $G(V,E)$ as follows:
Step 1: Variable Consistency Gadgets
For each variable $x_{i}$:
  • Create two vertices $v_{i}$ (representing $x_{i}$) and $\bar{v}_{i}$ (representing $\neg x_{i}$).
  • Connect $v_{i}$ and $\bar{v}_{i}$ with an edge of weight $W=3m+1$. This enforces consistency: in any optimal cut, $v_{i}$ and $\bar{v}_{i}$ must lie in opposite partitions.
Step 2: Clause Gadgets
For each clause $C_{j}=(\ell_{1}\lor\ell_{2}\lor\dots\lor\ell_{k})$:
  • Introduce two auxiliary vertices $u_{j}$ and $w_{j}$.
  • Connect $u_{j}$ to each literal vertex in $C_{j}$ (i.e., $v_{i}$ if $\ell=x_{i}$, or $\bar{v}_{i}$ if =$\ell=\neg x_{i}$) with edges of weight 1.
  • Connect $w_{j}$ to each literal vertex in $C_{j}$ with edges of weight 1.
  • Connect $u_{j}$ and $w_{j}$ with an edge of weight 1.
Step 3: Interpretation of Weights
  • The large weight $W=3m+1$ ensures that violating consistency (placing $v_{i}$ and $\bar{v}_{i}$ in the same partition) incurs a penalty exceeding any possible gain from clauses. Thus, optimal cuts correspond to valid truth assignments.
  • For clauses, the gadget ensures:
    • If at least one literal in $C_{j}$ is true (i.e., its vertex is in the opposite partition from $u_{j}$), the edges $(u_{j},\text{literal})$ and $(w_{j},\text{literal})$ contribute to the cut. This yields a total contribution of $3$ for the gadget (e.g., if $u_{j}$ is in $V_{1}$, edges from $u_{j}$ to true literals in $V_{2}$ and the edge $(u_{j},w_{j})$ contribute 3).
    • If all literals in $C_{j}$ are false, the gadget contributes only $2$ to the cut (edges $(u_{j},w_{j})$ and one literal edge).
Step 4: Correctness
Let $\text{OPT}_{\text{SAT}}$ and $\text{OPT}_{\text{Cut}}$ denote the optimal values of the Max-k-SAT and MaxCut instances, respectively. By construction:
$\text{OPT}_{\text{Cut}}=3\cdot\text{OPT}_{\text{SAT}}+nW.$
Since $W=3m+1$, the term $nW$ dominates, ensuring that any optimal cut must split all $(v_{i},\bar{v}_{i})$ pairs. The clause gadget contributions then directly encode satisfied clauses.
Step 5: Complexity
The graph $G$ has:
  • $2n+2m$ vertices (2 per variable, 2 per clause).
  • $O(n+km)$ edges (1 edge per variable, $12k+$ 1 edges per clause).
The construction is polynomial in $n$ and $m$. ∎
A.2 Reduction from MaxCut to Max-k-SAT
Theorem 2.
Any instance of MaxCut can be transformed into an equivalent instance of Max-k-SAT in polynomial time, preserving optimization structure.
Proof.
Given a MaxCut instance on a graph $G(V,E)$, construct a Max-2-SAT instance as follows:
Step 1: Variable Representation
For each vertex $v\in V$, create a Boolean variable $x_{v}$. The assignment $x_{v}=1$ (resp. $x_{v}=0$) encodes placing $v$ in partition $V_{1}$ (resp. $V_{2}$).
Step 2: Clause Construction
For each edge $(u,v)\in E$, add two clauses:
$(x_{u}\lor\neg x_{v})\quad\text{and}\quad(\neg x_{u}\lor x_{v}).$
These clauses are satisfied if and only if $u$ and $v$ are in different partitions.
Step 3: Correctness
The number of satisfied clauses equals twice the size of the cut. Thus, maximizing satisfied clauses corresponds to maximizing the cut value.
Step 4: Complexity
The reduction generates $2|E|$ clauses, which is polynomial in $|E|$. ∎
A.3 Conclusion
The reductions establish that Max-k-SAT and MaxCut are polynomial-time equivalent. This equivalence provides a theoretical framework to justify extending empirical result, i.e., the logistic saturation conjecture, between the two problems.
Appendix B Graphs used in the main text